# Coupling Generative Modeling and an Autoencoder with the Causal Bridge

### v1

@@ 7.1 Synthetic data
We first demonstrate the performance of the proposed method on two synthetic datasets introduced in (gretton_causal, ), and we perform the same experiments as considered there.
@@
Data The first experiment considers the Demand data introduced by (DeepIV, ). Details of the data generation process are found in Appendix E.1. The second experiment considers the dSprite data introduced by (matthey2020dsprites, ). Details of the data generation process can be found in Appendix E.2. For both datasets, we consider two sample size settings, 1000 and 5000, for a single dataset $\{(x,z,w,y)\}$, consistent with the setup in (gretton_causal, ).
@@
Metrics We estimate the average causal effect using the bridge function with (14) and compare it to the ground-truth causal effect in terms of the mean squared error (MSE). Note that this is only possible with synthetic data since the datasets used to train the model only have access to specific (factual) combinations of $\{(x,z,y)\}$, i.e., we do not also have values of $\{(z,y)\}$ for which the treatment $x$ takes values different (counterfactual) than those observed. The test points are evenly distributed over the treatment variable $X$ when calculating the out-of-sample MSE. Additional details are provided in Appendix E. This setup replicates the experiments introduced by (gretton_causal, ).
@@
Models considered We compare our method to the deep feature proxy variable (DFPV) method of (gretton_causal, ), which is considered the prior state of the art. In the results below, we consider the following model configurations: $i$) The original DFPV method gretton_causal . $ii$) The same underlying DFPV model (and hyperparameters), but replace iterative learning gretton_causal by first estimating $p(W|x,z)$, from which we sample when learning the bridge function. This setup allows examination of the benefits of learning to sample from $p(W|x,z)$, with no change to the form of $b(W,x)$ from DFPV. iiiiii) The causal bridge (CB) model learned with (12) and using the $p(W|x,z)$ sampling model. $iv$) The combined bridge and autoencoder (CB + AE) model with (12) and (13), i.e., $\mathcal{L}_{\theta_{Y}}+\mathcal{L}_{\theta_{X}}+\mathcal{L}_{\theta_{Z}}$. Concerning the new methods discussed in Section 6 (the last two methods above), our model architectures are relatively simple and not considerably more complex than those used by DFPV gretton_causal . This is done to demonstrate that the performance of our model variants is not attributed to overly complicated architectures. The details of the neural networks and hyperparameter tuning for all models are provided in Appendix F. Concerning the last two models discussed above: CB learned based on modeling $Y$ alone and CB + AE is based on the combined loss for modeling $(Y,X,Z$), the exact same models and hyperparameters are used for $g_{\theta_{Y}}(x,W,h_{\theta_{U}}(W,x,\epsilon))$ (this allows consideration of the impact of the autoencoder, without anything else changed). We also note that model selection for all approaches is based only on the loss for $Y$ in 12.
@@
Results Figure 3 shows that for experiments with synthetic data, Demand and dSprites, using our generator for $p(W|x,z)$ significantly improves performance relative to DFPV gretton_causal , which is likely due to the advantages of sampling (we draw 100 samples from $p(W|x,z)$ in these experiments). Consequently, we have the flexibility to generate many samples of $W$ for each combination of $\{(X=x,Z=z)\}$, enabling better learning and estimation of causal effects. Moreover, utilizing $h_{\theta_{U}}(W,x,\epsilon)$ in the generalized bridge model $g_{\theta_{Y}}(x,W,h_{\theta_{U}}(W,x,\epsilon))$, and thus moving beyond the DFPV representation of the bridge, further improves performance. The gains of this approach using (12) are most noticeable on the Demand dataset. Finally, incorporating the autoencoder for $(X,Z)$ using both (12) and (13) aids with the learning of the shared $h_{\theta_{U}}(W,x,\epsilon)$, particularly when the sample size is small, which is of particular interest in real-world scenarios.
@@ 7.2 Real-World data
Data Framingham is an observational longitudinal study designed to learn about the incidence and prevalence of cardiovascular disease (CVD) and its risk factors (benjamin1994independent, ). The data used here are the Offspring cohort, which consists of 3435 subjects split into 2404, 516, and 515 training, validation, and test samples, respectively. For this experiment, we are interested in estimating the average causal effect of taking statins (a cholesterol-lowing medication), i.e., the treatment $X=\{0,1\}$, on the timing $Y$ of future CVD events. These data are in the public domain, and therefore these experiments can be replicated.
Constructing proxies for Framingham The 32 covariates available for this dataset are not split in terms of treatment and outcome controls, $Z$ and $W$, respectively. Consequently, we adopt the proxy bucketing strategy of (TchetgenTchetgen2024AnIT, ), which divides covariates into $Z$ and $W$ according to their association with treatments and outcomes using effect sizes estimated from linear models on the training data. Details are provided in Appendix E.3.
Models considered We compare the proposed framework with three strong baselines: variants of balancing weighting schemes for the Cox proportional hazards (CoxPH) model for survival analysis rosenbaum1983central . Balancing weights are obtained from a linear logistic regression model built to estimate propensity scores $s=P(X=1|W=w,Z=z)$. Then, these weights are used to fit a CoxPH model where the input is only the treatment and the objective uses the balancing weights. We consider three weighting schemes: (CoxPH-Uniform) where all weights are 1; (CoxPH-IPW) using inverse probability weighting, $\omega=x/s+(1-x)/(1-s)$ cao2009improving ; (CoxPH-OW) using overlapping weighting, $\omega=x(1-s)+(1-x)s$ li2018balancing . We also consider the two proposed variants of the causal bridge, i.e., CB and CB + AE. Additional details about the CoxPH model and overlapping weights are provided in Appendix G. The sampler $p(W|x,z)$, the bridge and autoencoder architectures, and model selection details are provided in Appendix F.
Randomized control trial We also report the HR estimated from a separate randomized control trial (RCT) conducted specifically to assess the effects of statins on CVD outcomes yusuf2016cholesterol . These RCT results provide a powerful reference against which causal estimates can be compared.
@@ Appendix F Network Structures and Hyperparameters
In both synthetic data experiments, we use $k$-fold cross-validation ($k=5$) to select the learning rate from the range $10^{-3},10^{-4},10^{-5}]$, keeping the weight decay fixed at $10^{-5}$ without fine-tuning. The learning rate is the same for all model components, but one can try to make the learning rate different across different components. We can also use $k$-fold cross-validation for synthetic data and the validation split for Framingham data to determine the weights for each autoencoder loss accordingly. Once the hyperparameters are selected, the model is trained with a train/validation split of $0.9/0.1$. Early stopping is implemented using the validation loss of the causal bridge $\mathcal{L}_{\theta_{Y}}$ only, even for the autoencoder version of the model. In Framingham experiments, we use the validation split provided in the dataset to measure the concordance index (C-Index) and select the batch size from the range $[384,512,768]$, keeping the learning rate and weight decay fixed at $10^{-1}$ and $10^{-4}$, respectively. The test split in Framingham is used only to evaluate the C-Index of the trained model. To ensure a fair comparison, we utilize the training and validation split to learn the baseline CoxPH models with three different weighting schemes.
@@ F.1 Demand Experiments
For the dSprite experiment, we use the AdamW optimizer with a weight decay of $10^{-5}$ and a learning rate of $10^{-4}$ for all models. The autoencoder loss $\mathcal{L}_{\theta_{X}}$ is weighted with $\mathtt{w_{x}}=0.01$, and $\mathcal{L}_{\theta_{Z}}$ is weighted with $\mathtt{w_{z}}=0.01$.
@@ 3 FC(4096,1)(4096,1), Sigmoid
2 FC(8692,256)(8692,256), ReLU
3 FC(256,32)(256,32)

### v2

@@ 7.1 Synthetic data
We first demonstrate the performance of the proposed method on two synthetic datasets introduced in (Xu et al., 2021b), and we perform the same experiments as considered there.
@@
Data The first experiment considers the Demand data introduced by (Hartford et al., 2017). Details of the data generation process are found in Appendix E.1. The second experiment considers the dSprite data introduced by (Matthey et al., 2020). Details of the data generation process can be found in Appendix E.2. For both datasets, we consider two sample size settings, 1000 and 5000, for a single dataset $\{(x,z,w,y)\}$, consistent with the setup in (Xu et al., 2021b).
@@
Metrics We estimate the average causal effect using the bridge function with (14) and compare it to the ground-truth causal effect in terms of the mean squared error (MSE). Note that this is only possible with synthetic data since the datasets used to train the model only have access to specific (factual) combinations of $\{(x,z,y)\}$, i.e., we do not also have values of $\{(z,y)\}$ for which the treatment $x$ takes values different (counterfactual) than those observed. The test points are evenly distributed over the treatment variable $X$ when calculating the out-of-sample MSE. Additional details are provided in Appendix E. This setup replicates the experiments introduced by (Xu et al., 2021b).
@@
Models considered We compare our method to the deep feature proxy variable (DFPV) method of (Xu et al., 2021b), which is considered the prior state of the art. In the results below, we consider the following model configurations: $i$) The original DFPV method Xu et al. (2021b). $ii$) The same underlying DFPV model (and hyperparameters), but replace iterative learning Xu et al. (2021b) by first estimating $p(W|x,z)$, from which we sample when learning the bridge function. This setup allows examination of the benefits of learning to sample from $p(W|x,z)$, with no change to the form of $b(W,x)$ from DFPV. iiiiii) The causal bridge (CB) model learned with (12) and using the $p(W|x,z)$ sampling model. $iv$) The combined bridge and autoencoder (CB + AE) model with (12) and (13), i.e., $\mathcal{L}_{\theta_{Y}}+\mathcal{L}_{\theta_{X}}+\mathcal{L}_{\theta_{Z}}$. Concerning the new methods discussed in Section 6 (the last two methods above), our model architectures are relatively simple and not considerably more complex than those used by DFPV Xu et al. (2021b). This is done to demonstrate that the performance of our model variants is not attributed to overly complicated architectures. The details of the neural networks and hyperparameter tuning for all models are provided in Appendix F. Concerning the last two models discussed above: CB learned based on modeling $Y$ alone and CB + AE is based on the combined loss for modeling $(Y,X,Z$), the exact same models and hyperparameters are used for $g_{\theta_{Y}}(x,W,h_{\theta_{U}}(W,x,\epsilon))$ (this allows consideration of the impact of the autoencoder, without anything else changed). We also note that model selection for all approaches is based only on the loss for $Y$ in 12.
@@
Results Figure 3 shows that for experiments with synthetic data, Demand and dSprites, using our generator for $p(W|x,z)$ significantly improves performance relative to DFPV Xu et al. (2021b), which is likely due to the advantages of sampling (we draw 100 samples from $p(W|x,z)$ in these experiments). Consequently, we have the flexibility to generate many samples of $W$ for each combination of $\{(X=x,Z=z)\}$, enabling better learning and estimation of causal effects. Moreover, utilizing $h_{\theta_{U}}(W,x,\epsilon)$ in the generalized bridge model $g_{\theta_{Y}}(x,W,h_{\theta_{U}}(W,x,\epsilon))$, and thus moving beyond the DFPV representation of the bridge, further improves performance. The gains of this approach using (12) are most noticeable on the Demand dataset. Finally, incorporating the autoencoder for $(X,Z)$ using both (12) and (13) aids with the learning of the shared $h_{\theta_{U}}(W,x,\epsilon)$, particularly when the sample size is small, which is of particular interest in real-world scenarios.
@@ 7.2 Real-World data
Data Framingham is an observational longitudinal study designed to learn about the incidence and prevalence of cardiovascular disease (CVD) and its risk factors (Benjamin et al., 1994). The data used here are the Offspring cohort, which consists of 3435 subjects split into 2404, 516, and 515 training, validation, and test samples, respectively. For this experiment, we are interested in estimating the average causal effect of taking statins (a cholesterol-lowing medication), i.e., the treatment $X=\{0,1\}$, on the timing $Y$ of future CVD events. These data are in the public domain, and therefore these experiments can be replicated.
Constructing proxies for Framingham The 32 covariates available for this dataset are not split in terms of treatment and outcome controls, $Z$ and $W$, respectively. Consequently, we adopt the proxy bucketing strategy of (Tchetgen et al., 2024), which divides covariates into $Z$ and $W$ according to their association with treatments and outcomes using effect sizes estimated from linear models on the training data. Details are provided in Appendix E.3.
Models considered We compare the proposed framework with three strong baselines: variants of balancing weighting schemes for the Cox proportional hazards (CoxPH) model for survival analysis Rosenbaum and Rubin (1983b). Balancing weights are obtained from a linear logistic regression model built to estimate propensity scores $s=P(X=1|W=w,Z=z)$. Then, these weights are used to fit a CoxPH model where the input is only the treatment and the objective uses the balancing weights. We consider three weighting schemes: (CoxPH-Uniform) where all weights are 1; (CoxPH-IPW) using inverse probability weighting, $\omega=x/s+(1-x)/(1-s)$ Cao et al. (2009); (CoxPH-OW) using overlapping weighting, $\omega=x(1-s)+(1-x)s$ Li et al. (2018). We also consider the two proposed variants of the causal bridge, i.e., CB and CB + AE. Additional details about the CoxPH model and overlapping weights are provided in Appendix G. The sampler $p(W|x,z)$, the bridge and autoencoder architectures, and model selection details are provided in Appendix F.
Randomized control trial We also report the HR estimated from a separate randomized control trial (RCT) conducted specifically to assess the effects of statins on CVD outcomes Yusuf et al. (2016). These RCT results provide a powerful reference against which causal estimates can be compared.
@@ Appendix F Network Structures and Hyperparameters
In both synthetic data experiments, we use $k$-fold cross-validation ($k=5$) to select the learning rate from the range $10^{-3},10^{-4},10^{-5}]$. The learning rate is the same for all model components, but one can try to make the learning rate different across different components. We can also use $k$-fold cross-validation for synthetic data and the validation split for Framingham data to determine the weights for each autoencoder loss accordingly. Once the hyperparameters are selected, the model is trained with a train/validation split of $0.8/0.2$. Early stopping is implemented using the validation loss of the causal bridge $\mathcal{L}_{\theta_{Y}}$ only, even for the autoencoder version of the model. In Framingham experiments, we use the validation split provided in the dataset to measure the concordance index (C-Index) and select the batch size from the range $[384,512,768]$, keeping the learning rate and weight decay fixed at $10^{-1}$ and $10^{-4}$, respectively. The test split in Framingham is used only to evaluate the C-Index of the trained model. To ensure a fair comparison, we utilize the training and validation split to learn the baseline CoxPH models with three different weighting schemes.
@@ F.1 Demand Experiments
For the dSprite experiment, we use the Adam with learning rate of $10^{-3}$ for all models. The autoencoder loss $\mathcal{L}_{\theta_{X}}$ is weighted with $\mathtt{w_{x}}=100$, and $\mathcal{L}_{\theta_{Z}}$ is weighted with $\mathtt{w_{z}}=0.01$.
@@ 3 FC(4096,1)(4096,1), Sigmoid
2 FC(8692,1024)(8692,1024), ReLU
3 FC(1024,256)(1024,256), ReLU
4 FC(256,64)(256,64), ReLU
5 FC(64,32)(64,32)
Model $\phi_{X}$
1 Input(x)(x)
2 FC(4096,64)(4096,64), ReLU
3 FC(64,32)(64,32)